# Potato Potahto in the FAO-GAEZ Productivity Measures?

### v2

@@ Appendix B Simple Examples
where $\tilde{Z}\coloneqq\frac{Z_{1,res}+\tilde{V}}{2}$ and $\tilde{V}$ is a random variable satisfying $Cov\left(V_{1},Y\right)=0$, $Cov\left(Z_{1,res},V_{1}\right)=0$ and $Var\left(V_{1}\right)=1$. In other words, $\tilde{Z}$ is the average of $Z_{1,res}$ and a noise term that has the same variance of $X_{res}^{*}$.
@@ Appendix D No Prior Knowledge of the Target Parameter’s Sign
The reduced-form coefficients identify the sign of the true target parameter, i.e., $sign\left\{Cov\left(Z_{1},Y_{res}\right)\right\}=sign\left\{\left(Z_{2},Y_{res}\right)\right\}=sign\left\{\beta\right\}$. [^14] The $sign$ function equals one if its argument is a positive number, zero if its argument is zero, and minus one if its argument is a strictly negative number.
@@ Appendix F Applications Using Other Datasets
In this appendix, we discuss other empirical contexts where our partial identification strategy may be useful. To illustrate the broad applicability of our methodology, we will briefly analyze three empirical examples where the variable of interest is measured with error, and there might exist two possible proxies.
First, Gali and Gertler (1999) aims to estimate a New Keynesian Phillips Curve. To do so, they need to measure real marginal costs. However, aggregate marginal costs are not found in standard national accounts. To circumvent this issue, they propose to use a model-based proxy for marginal cost. In their Cobb-Douglas production model, the unobservable marginal cost is equal to the labor income share. However, different production models may suggest different proxies for the marginal cost variable. In this scenario, there would be more than one possible measurement of the variable of interest and our partial identification strategy may be used to estimate a New Keynesian Phillips Curve.
@@
Second, many recent studies (La Ferrara et al., 2012; Adena et al., 2015) focus on the impact of media access on economic behavior. To do so, they frequently measure media access by combining an irregular terrain model and the location of TV or radio transmitters. Since the most commonly used irregular terrain model is in its third version (Hufford, 1995), this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5.
@@
Third, Rodrik and Wacziarg (2005) analyze the impact of democratic transitions on economic outcomes. To measure their treatment variable, they use the Polity IV index. This expert-coded measure of democracy levels is in its fourth version and is based on experts’ interpretation of political regimes in each country and each year. Since different experts might interpret the same political regime differently, there is measurement error in each version of this index. Although this setting is similar to the FAO-GAEZ productivity measures (Sections 2.1 and 5), we believe that Assumption 5 may not be appropriate in this context. Since political experts might overreact to reanalyses of historical phenomena, it may be the case that the measurement error terms in different versions of the Polity Index are negatively correlated. In this case, the bounds proposed in Appendix C represent a more appropriate approach than the bounds in the main text.

### v3

@@ Appendix B Simple Examples
where $\tilde{Z}\coloneqq\frac{Z_{1,res}+\tilde{V}}{2}$ and $\tilde{V}$ is a random variable satisfying $Cov\left(\tilde{V},Y\right)=0$, $Cov\left(Z_{1,res},\tilde{V}\right)=0$ and $Var\left(\tilde{V}\right)=1$. In other words, $\tilde{Z}$ is the average of $Z_{1,res}$ and a noise term that has the same variance of $X_{res}^{*}$.
@@ Appendix D No Prior Knowledge of the Target Parameter’s Sign
The reduced-form coefficients identify the sign of the true target parameter, i.e., $sign\left\{Cov\left(Z_{1},Y_{res}\right)\right\}=sign\left\{Cov\left(Z_{2},Y_{res}\right)\right\}=sign\left\{\beta\right\}$. [^15] The $sign$ function equals one if its argument is a positive number, zero if its argument is zero, and minus one if its argument is a strictly negative number.
@@ Appendix F Applications Using Other Datasets
In this appendix, we discuss other empirical contexts where our partial identification strategy may be useful. To illustrate the broad applicability of our methodology, we will briefly analyze six empirical examples where the variable of interest is measured with error, and there might exist two possible proxies.
@@
First, many recent studies (Olken, 2009; Adena et al., 2015) focus on the impact of media access on economic behavior. To do so, they frequently measure media access by combining an irregular terrain model and the location of TV or radio transmitters. Since the most commonly used irregular terrain model is in its third version (Hufford, 1995), this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5.
Second, many articles (Post and Hanewald, 2013; Gan et al., 2015; Bissonnette et al., 2017) discuss the impact of subjective beliefs about life expectancy on economic behavior (e.g., savings, investment, bequests). To measure this type of belief, many surveys ask individuals about their subjective probability of attaining various ages. However, individual responses frequently present measurement issues, such as rounding and focal answers. To address these problems, many authors adopt different imputation methods that approximate the true subjective probability of attaining a certain age. By treating the results of each imputation model as a different proxy, our proposed method may be used in this empirical setting.
Third, multiple studies (Hsieh and Klenow, 2009, 2010) are interested in the impact of total factor productivity on economic outcomes, such as plant exit and income levels. To measure total factor productivity, different authors adopt different models (Olley and Pakes, 1996; Hsieh and Klenow, 2009; Ackerberg et al., 2015). By treating the measures produced by each model as a different proxy for the true latent total factor productivity, our partial identification strategy may be used in this empirical context.
Fourth, economic historians (Sala-i-Martin, 1996; Hartley, 2021) frequently focus on understanding GDP growth and fluctuations in the long run. To do so, they need measures of income levels over long periods of time, including moments when national accounts were not collected in a systematic way. To access this type of variable, most authors use the Maddison Project Data (Bolt and van Zanden, 2025), which is in its fifth version. Each version varies either because new countries were included in the sample or, more importantly to us, new information was collected about existing countries and their historical income data were updated. Since different versions of these variables may have different measurement error terms when compared to the true income level of each country, this setting is very similar to the FAO-GAEZ productivity measures discussed in Sections 2.1 and 5.
@@
Fifth, Rodrik and Wacziarg (2005) analyze the impact of democratic transitions on economic outcomes. To measure their treatment variable, they use the Polity IV index. This expert-coded measure of democracy levels is in its fourth version and is based on experts’ interpretation of political regimes in each country and each year. Since different experts might interpret the same political regime differently, there is measurement error in each version of this index. Although this setting is similar to the FAO-GAEZ productivity measures (Sections 2.1 and 5), we believe that Assumption 5 may not be appropriate in this context. Since political experts might overreact to reanalyses of historical phenomena, it may be the case that the measurement error terms in different versions of the Polity Index are negatively correlated. In this case, the bounds proposed in Appendix C represent a more appropriate approach than the bounds in the main text.
Sixth, multiple articles (Alfaro et al., 2008; Rajkumar and Swaroop, 2008; Aksoy et al., 2025) are interested in the impact of a country’s governance quality on economic outcomes, such as health quality, political trust and international capital flows. To measure governance quality, Alfaro et al. (2008) and Rajkumar and Swaroop (2008) use a variety of measures from the Political Risk Services Group, while Aksoy et al. (2025) uses the V-DEM index and one measure from the Worldwide Governance Indicators. By treating these different measures as different proxies for the true unobservable governance quality, our partial identification strategy may be applied in this empirical context.